OpenStudy (anonymous):

Can someone please help me quickly? Write the equation of the line that passes through (2, 2) and (6, 3) in standard form. I know you have to get the slope, and the slope is 1/4. But then you have to do something else and that's where I'm stuck.. I have no idea what I'm supposed to do now...

OpenStudy (zpupster):

y—y1=m(x—x1) (2,2) plug them in y-2=1/4(x-2) or you can plug in the other ordered pair (6,3) y-3=1/4(x-6) if you like --you can solve y and put in this form y=mx+b

OpenStudy (anonymous):

How do I solve y then, because I can't have it in that format.. I need it in y=mx+b ;-;

OpenStudy (zpupster):

y-2=1/4(x-2) first mult each term inside pareetheses by 1/4 on RHS y-2 = 1/4x-1/2 add 2 to both sides y= 1/4x-1/2+4/2 y= 1/4x+3/2

OpenStudy (texaschic101):

(3 - 2) / (6 - 2) = 1/4 -- your correct with the slope now we can either solve it by using y = mx + b and solve for b or we can use y - y1 = m(x - x1) lets do y = mx + b slope(m) = 1/4 (2,2) x = 2 and y = 2 now we sub 2 = 1/4(2) + b 2 = 1/2 + b 2 - 1/2 = b 4/2 - 1/2 = b 3/2 = b therefore your line is : y = 1/4x + 3/2 Now we need standard form... y = 1/4x + 3/2 (multiply the equation by 4) 4y = x + 6 (now subtract x) -x + 4y = 6 (now multiply by -1 to make x positive) x - 4y = - 6 or y - y1 = m(x - x1) slope(m) = 1/4 (2,2) x1 = 2 and y1 = 2 now we sub y - 2 = 1/4(x - 2) y - 2 = 1/4x - 1/2 y = 1/4x - 1/2 + 2 y = 1/4x - 1/2 + 4/2 y = 1/4x + 3/2 (subtract 1/4x from both sides) -1/4x + y = 3/2 (multiply by 4) -x + 4y = 6 (multiply by -1) x - 4y = -6 you will arrive at the same answer either way you do it

OpenStudy (anonymous):

Thank you so much Texas! :DD I can base myself off of that and finish my other questions now. c: Thank you again!

OpenStudy (texaschic101):

anytime :) glad to help